# Large Language Models are Contrastive Reasoners

### v1

@@ 4.2 Results
Table 1 presents a summary of the accuracy achieved by our method (Zero-shot-CP and Zero-shot-CoT-CP), standard zero-shot prompting (Zero-shot), and Zero-shot-CoT across different datasets. Notably, Zero-shot-CP demonstrates significant improvements over Zero-shot on all 12 datasets across various tasks using gpt-35-turbo. For instance, Zero-shot-CP achieves score gains ranging from $14.3\%$ to $73.2\%$ on GSM8K, from $61.2\%$ to $95.2\%$ on MultiArith and from $4.2\%$ to $41.8\%$ on Last Letter Concatenation. Moreover, Zero-shot-CP outperforms Zero-shot on the majority (9 out of 12) of datasets when using GPT-4, with improvements ranging from $35.9\%$ to $88.8\%$ on GSM8K and from $41.3\%$ to $62.2\%$ on AQUA-RAT. These results indicate that eliciting self-awareness in LLMs to compare incorrect and correct answers can help prevent incorrect responses.
Zero-shot-CP outperforms Zero-shot-CoT in the majority (4 out of 6) of arithmetic reasoning tasks, suggesting that the self-awareness of LLMs regarding incorrect answers may be more crucial than their self-awareness regarding steps in mathematical reasoning. Furthermore, in commonsense reasoning tasks, Zero-shot-CP consistently outperforms Zero-shot, while Zero-shot-CoT exhibits inferior results. This is likely because commonsense reasoning tasks require fewer steps, making awareness of individual pieces of commonsense knowledge more crucial. However, Zero-shot CoT performs worse than Zero-shot-CP in symbolic reasoning and other reasoning tasks, indicating that for tasks like Coin Flip and Shuffled Objects, the steps involved are more crucial. Moreover, breaking down the reasoning process into individual steps makes it less likely for LLMs to make mistakes in these tasks, as the action space is very limited. Nevertheless, by combining Zero-shot-CP and Zero-shot-CoT, Zero-shot-CoT-CP achieves better results in most tasks.
@@
Zero-shot-CoT-CP A: Let’s think step by step and give both a correct answer and a wrong answer. Correct Answer:
       Danny found 50 bottle caps and 46 wrappers at the park. So, he found 50 - 46 = 4 more bottle caps
       than wrappers at the park.
       Incorrect Answer:
       If we mistakenly subtract the number of bottle caps and wrappers in his collection from the number
       he found at the park, we would get: 50 - 21 = 29 bottle caps and 46 - 52 = -6 wrappers. This would
       incorrectly suggest that Danny found 29 - (-6) = 35 more bottle caps than wrappers at the park.
       Therefore, the correct answer (arabic numerals) is 4
@@
Table 3: An example question from CommonsenseQA and responses from Zero-shot, Zero-shot-CoT, Zero-shot-CP and Zero-shot-CoT-CP. We use GPT-4 as the base model.
@@
Zero-shot-CoT-CP A: Let’s think step by step and give both a correct answer and a wrong answer. Correct Answer: (D) painting.
       A bald eagle can be depicted in a painting, as it is a common symbol of strength and freedom,
       especially in American culture.
       Incorrect Answer: (A) rural area. While it’s true that bald eagles can be found in rural areas,
       the phrasing of the question suggests a type of work or creation where a bald eagle might be represented,
       not its natural habitat.
       Therefore, among A through E, the correct answer is (D) painting.
@@
Qualitative Analysis
Table 2 and Table 3 illustrates examples from SVAMP and CommonsenseQA. For the example from the arithmetic reasoning task SVAMP, we found that the reasoning process of zero-shot is correct, but it produces an incorrect answer "-4". Zero-shot-CoT is disrupted by irrelevant information, resulting in incorrect reasoning processes and answers being generated. Zero-shot-CP, on the other hand, is not disrupted and provides both the correct answer and explanation. Combining Zero-shot-CoT with contrastive prompting also yields correct answers. For the example from the common sense reasoning task CommonsenseQA, contrastive prompting is able to recognize the word "work" in the question and provide the correct answer, while Zero-shot and Zero-shot-CoT cannot.
In Appendix B, we present responses generated by Zero-shot-CP for each dataset. Figure 415 gives both a positive example and a negative example of Zero-shot-CP on each dataset. From positive examples, we found that Zero-shot-CP can generate "wrong" answers that are indeed incorrect in most cases (11/12), except for Tracking Shuffled Object (Figure 12). Incorrect answers are generated by intentionally calculating inaccurately (Figure 11), disregarding important details (Figure 9), searching for descriptions that are not present in the question (Figure 8), or deliberately providing explanations that contradict common sense (Figure 14). From negative examples, We found that the "wrong answers" provided by Zero-shot-CP can actually be valid answers (Figure 5, 6, 7, 11, 13 and 14). In some other negative examples, both the "correct answers" and "incorrect answers" provided by Zero-shot-CP are inconsistent with the ground truth (Figure 4, 8, 9, 10 and 15). Furthermore, we manually annotated 10 solved problems and 10 unsolved problems of Zero-shot-CP for each of the 12 datasets. Table 4 provides the categorization and counts of these 120 solved problems and 120 unsolved problems. We found that for the solved problems, the majority (112/120) of the given "wrong" answers were indeed incorrect. For the unsolved problems, the majority (91/120) of both the "correct" and "wrong" answers were incorrect, with a portion (23/120) of the "wrong" answers actually being the ground truth. This situation typically occurs in yes or no questions.
Category # Examples
The given "correct" answer is the GT, no "wrong" answer is given. 4
The given "correct" answer is incorrect, no "wrong" answer is given. 6
Table 5: Comparison prompting templates using accuracies (in percentage) on AQUA-RAT, GSM8K, AddSub and MultiArith in zero-shot setting. gpt-35-turbo is used as the model. Bolded numbers indicate the best results within each block’s column, while underlined numbers indicate the best results across the entire column.
AQUA GSM8K AddSub MultiArith
Let’s give a correct and a wrong answer. 40.2 73.2 90.6 95.2
Let’s first give a wrong answer, then give the correct answer. 53.5 71.0 84.3 90.7
Let’s first give the correct answer, then give a wrong answer. 35.8 72.6 91.1 94.8
Let’s give a correct and an incorrect answer. 40.6 72.5 90.4 94.2
Please give a correct and a wrong answer. 35.8 69.2 88.9 90.5
Let’s think step by step and give both a correct answer and a wrong answer. 60.6 73.5 88.6 96.2
Let’s give a correct and a wrong answer. Let’s also think 32.7 71.9 89.1 95.5
step by step for the correct and the wrong answer.
Zero-shot 29.9 14.3 82.8 61.2
Zero-shot-COT (Let’s think step by step.) 55.9 75.1 86.6 94.8
The impact of prompt selection on Zero-shot-CP
We explore different contrastive prompts and their combination with Zero-shot-CoT. Table 5 outlines performance using 9 different templates with three classes. The first category is contrastive prompts only, we found "Let’s give a correct and a wrong answer." achieves the best results. "Let’s first give a wrong answer, then give the correct answer." performs well on AQUA-RAT but it performs worse on other datasets. "Let’s first give the correct answer, then give a wrong answer." generally performs well on the four datasets, meaning that providing the correct answer first and then the incorrect answer generally leads to better results. The trigger word "incorrect" performs slightly worse than "wrong", and the trigger word "Please" performs much worse than "Let’s". This is likely because, in the pre-training and fine-tuning data, there are slightly fewer occurrences of "incorrect" compared to "wrong" in samples related to correct and incorrect answers, and "Please" is rarely present as this type of data is generally not dialogue data. The second type of template combines CP with Zero-shot-CoT, and we found that starting with the steps performs better than starting with the correct and wrong answers. The third type of template is the standard Zero-shot and Zero-shot-CoT, which we have compared before. Overall, it appears that Zero-shot-CoT-CP ("Let’s think step by step and give both a correct answer and a wrong answer.") performs the best.
The impact of number of wrong answers on Zero-shot-CP
We explored the impact of the number of incorrect answers on accuracy. We vary the number of wrong answers from 0 to $4$, where 0 means standard zero-shot prompting. For $k=1,2,3,4$, we use the template "Let’s give a correct and $k^{\prime}$ wrong answer(s).", where $k^{\prime}$ can be "a", "two", "three" and "four". Figure 3 plots the results. We found that providing 1-2 incorrect answers yielded the best results in general. The only exception is on AQUA-RAT, where providing more incorrect answers resulted in higher accuracy. This is because the task involves a multiple-choice question with five options, and excluding more incorrect answers makes the LLMs more certain about the correct answer. For mathematical reasoning tasks with an infinite number of answers, providing just one incorrect answer seems to be sufficient.
(a) AQUA
(b) GSM8K
(c) AddSub
(d) MultiArith
Figure 3: Accuracy scores by varying the number of wrong answers. We test GPT-4 and gpt-35-turbo on (a) AQUA-RAT, (b) GSM8K, (c) AddSub and (d) MultiArith. The range of the number of wrong answers is from 0 (Zero-shot) to 4.
Table 6: Comparison with baseline methods using accuracies (in percentage) on MultiArith, GSM8K, StrategyQA, AQUA-RAT and SVAMP. gpt-35-turbo is used as the model if not specified. The baseline results with citations are obtained from corresponding papers. Bolded numbers indicate the best results within each block’s column, while underlined numbers indicate the best results across the entire column.
MultiArith GSM8K StrategyQA AQUA SVAMP
Zero-shot 61.2 14.3 65.0 29.9 69.7
Zero-shot-CoT 94.8 75.1 60.9 55.9 81.9
Zero-shot-CP 95.2 73.2 67.3 40.2 83.2
Zero-shot-CoT-CP 96.2 73.5 66.7 60.6 85.9
Few-shot 87.3 58.2 56.7 37.4 78.2
Few-shot-CoT 98.0 71.1 62.2 55.5 81.0
Few-shot-CoT-CP 97.5 72.7 68.7 52.0 82.2
Few-shot-CoT (GPT-4) 98.3 89.5 79.1 58.7 83.3
Few-shot-CoT-CP (GPT-4) 98.7 90.3 78.2 66.9 91.8
Contrastive CoT [4] 79.0 66.2 57.5 81.6
Self-consistency (Code-davinci-002) [19] 100.0 78.0 79.8 52.0 86.8
PAL (Codex) [6] 99.2 80.4 79.4
Zero-shot-PoT (Codex) [3] 92.2 57.0 43.9 70.8
Few-shot-PoT (Codex) [3] 71.6 54.1 85.2
Few-shot-PoT-SC (Codex) [3] 80.0 58.6 89.1
ToT (GPT-4) [22] 90.0 83.0
GoT (T5-large) [23] 82.2
Self-generated Exemplars [24] 77.8
Self-Refine [11] 75.1
LEAP [25] 77.4
Zero-Shot-CoT + RCI  [8] 97.2 86.2 85.8
Few-Shot-CoT + RCI  [8] 99.2 84.3 87.4
Comparison with other baselines
Table 6 compares the performances on four mathematical reasoning datasets (MultiArith, GSM8K, AQUA-RAT and SVAMP) and one common sense reasoning dataset (StrategyQA) across CP and baselines. We find that Zero-shot-CP not only outperforms Few-shot, but also achieves comparable or even superior results to Few-shot-CoT. For instance, on GSM8K, the absolute accuracy has improved by $2.1\%$, and on StrategyQA, the absolute accuracy has improved by $5.1\%$. This suggests that in certain cases, the provided examples and reasoning steps may not be as effective as directly triggering the LLM’s self-awareness of errors. By combining CP and Few-shot-CoT, we can achieve even better results. Furthermore, if we utilize the GPT-4 model, we can attain performance that is comparable to or even superior to the current state-of-the-art methods. For example, in AQUA, SVAMP, and GSM8K, we have achieved higher accuracy scores compared to recently published results.
The main reasons why CP works well are threefold: 1) the pre-training data of LLMs is very likely to contain a variety of correct and incorrect answers to different types of questions. For instance, many web pages and books provide correct and incorrect answers to math reasoning [^2] https://prek-math-te.stanford.edu/operations/analyzing-thinking-underlying-wrong-answers [^3] https://mathmistakes.org/category/elementary-school/ [^4] https://www.gutenberg.org/ebooks/38769 and common sense reasoning   [^5] https://www.proprofs.com/quiz-school/story.php?title=common-sense-quiz_1666https://www.wikihow.com/Common-Sense-Quiz questions. Answers to questions on social media platforms like Reddit, Quora, and Zhihu can be voted on by others through “upvotes” or “downvotes”. Highly upvoted answers are more likely to be correct answers while others may be incorrect. Pre-training LLMs with massive text containing these correct and wrong answers can encode general patterns (token probability) of these answers in LLM parameters. When prompted with contrastive prompts, LLMs can leverage these patterns to generate both a correct and a wrong answer. The "correct" answer is more likely to align with ground truth, as the model has learned to eliminate possible wrong answers. 2) Instruction tuning unlocks the abilities of LLMs to give correct and incorrect answers by fine-tuning on various natural language processing tasks including reasoning tasks [20]. 3) RLHF fine-tunes LLMs using human feedback data, which offers relative judgments on the quality of answers. This feedback is valuable for enhancing the LLMs’ capability to distinguish between correct and incorrect answers.
5 Limitations and future directions
Our work has some limitations and there is room for further exploration and improvement. Firstly, we have not yet validated the effectiveness of contrastive prompting on smaller models such as LLaMA-13B, LLaMA-7B [18], and ChatGLM-6B [5]. Secondly, the generated incorrect answers may contain harmful content or biases, which need to be further mitigated. Thirdly, we can further explore the combination of contrastive prompting with other prompting methods, such as self-consistency. Lastly, exploring the impact of contrastive prompting on LLM parameters and visualizing it would be an interesting future direction.
6 Conclusion
We propose Contrastive Prompting (CP), a template-based prompting approach for contrastive reasoning. Quantitative and qualitative results indicate that Zero-Shot-CoT shows significant improvements across various reasoning tasks. Our method can seamlessly integrate with any prompting technique by incorporating a trigger sentence before the LLM provides answers. CP not only outperforms zero-shot CoT and few-shot CoT in the majority of arithmetic and commonsense reasoning tasks, but also achieves comparable or even superior results when compared to state-of-the-art methods.
Appendix A Details of Experimental Setup
A.1 Code, Prompts, Logs
All code is available at https://github.com/yao8839836/cp.
A.2 Prompts For Answer Extraction
Appendix B Additional Experimental Results
In this section, we provide a summary of additional example texts generated by Zero-shot-CP. gpt-35-turbo is used as the model if not specified. Figure 415 show a positive example and a negative example of Zero-shot-CP on each dataset. "GT" in the figures means "Ground Truth".
The 240 examples, along with our annotations, can be accessed at the following link: https://github.com/yao8839836/cp/blob/master/results/zero_shot_cp_gpt4_240_examples_labeled.txt.
Table 7: Answer extraction prompts used in our CP experiments.

### v3

@@ 4.2 Results
Table 1 presents the accuracy scores achieved by our Zero-shot-CP, standard zero-shot prompting (Zero-shot) and Zero-shot-CoT across five datasets. We ran all methods five times using GPT-3.5-Turbo and report the average scores. We found that the differences in each run were minimal. Zero-shot-CP consistently outperformed Zero-shot-CoT and Zero-shot across most (4 out of 5) datasets.
Table 5 in Appendix B presents more comprehensive results. Notably, Zero-shot-CP demonstrates significant improvements over Zero-shot on all 12 datasets across various tasks using GPT-3.5-Turbo. For instance, Zero-shot-CP achieves score gains ranging from $14.3\%$ to $73.2\%$ on GSM8K, from $61.2\%$ to $95.2\%$ on MultiArith and from $4.2\%$ to $41.8\%$ on Last Letter Concatenation. Moreover, Zero-shot-CP outperforms Zero-shot on the majority (9 out of 12) of datasets when using GPT-4, with improvements ranging from $35.9\%$ to $88.8\%$ on GSM8K and from $41.3\%$ to $62.2\%$ on AQUA-RAT. These results indicate that eliciting self-awareness in LLMs to compare incorrect and correct answers can help prevent incorrect responses.
Zero-shot-CP outperforms Zero-shot-CoT in the majority (4 out of 6) of arithmetic reasoning tasks, suggesting that the self-awareness of LLMs regarding incorrect answers may be more crucial than their self-awareness regarding steps in mathematical reasoning. Furthermore, in commonsense reasoning tasks, Zero-shot-CP consistently outperforms Zero-shot (2 out of 2), while Zero-shot-CoT exhibits inferior results. This is likely because commonsense reasoning tasks require fewer steps, making awareness of individual pieces of commonsense knowledge more crucial. However, Zero-shot-CP performs worse than Zero-shot-CoT in symbolic reasoning and other reasoning tasks, indicating that for tasks like Coin Flip and Shuffled Objects, the steps involved are more crucial. Moreover, breaking down the reasoning process into individual steps makes it less likely for LLMs to make mistakes in these tasks, as the action space is very limited. Nevertheless, by combining Zero-shot-CP and Zero-shot-CoT, Zero-shot-CoT-CP achieves better results in most tasks (see Table  4).
Table 8 in Appendix B presents the results of using various open-source LLMs: LLaMA3-8B, LLaMA3-70B, ChatGLM3-6B, and Qwen1.5-72B-Chat as base models. The results demonstrate that Zero-shot-CP not only performs well with state-of-the-art GPT models, but also exhibits significant improvements across multiple sizes of open-source models.
@@
Qualitative Analysis
Table 2 and Table 6 in Appendix  B illustrates examples from SVAMP and CommonsenseQA. For the example from the arithmetic reasoning task SVAMP, we found that the reasoning process of zero-shot is correct, but it produces an incorrect answer "-4". Zero-shot-CoT is disrupted by irrelevant information, resulting in incorrect reasoning processes and answers being generated. Zero-shot-CP, on the other hand, is not disrupted and provides both the correct answer and explanation. We can see that the "wrong answer" "-4" from Zero-shot-CP is a real mistake made by Zero-shot. For the example from the common sense reasoning task CommonsenseQA, contrastive prompting is able to recognize the word "work" in the question and provide the correct answer, while Zero-shot and Zero-shot-CoT cannot.
In Appendix B, we present responses generated by Zero-shot-CP for each dataset. Figure 516 gives both a positive example and a negative example of Zero-shot-CP on each dataset. From positive examples, we found that Zero-shot-CP can generate "wrong" answers that are indeed incorrect in most cases (11/12), except for Tracking Shuffled Object (Figure 13). Incorrect answers are generated by intentionally calculating inaccurately (Figure 12), disregarding important details (Figure 10), searching for descriptions that are not present in the question (Figure 9), or deliberately providing explanations that contradict common sense (Figure 15). From negative examples, We found that the "wrong answers" provided by Zero-shot-CP can actually be valid answers (Figure 6, 7, 8, 12, 14 and 15). In some other negative examples, both the "correct answers" and "incorrect answers" provided by Zero-shot-CP are inconsistent with the ground truth (Figure 5, 9, 10, 11 and 16). From the figures, we found that Zero-shot-CP also outputs reasoning steps in the process of generating correct and incorrect answers, especially for arithmetic reasoning tasks. Furthermore, we manually annotated 10 solved problems and 10 unsolved problems of Zero-shot-CP for each of the 12 datasets. Table 10 provides the categorization and counts of these 120 solved problems and 120 unsolved problems. We found that for the solved problems, the majority (112/120) of the given "wrong" answers were indeed incorrect. For the unsolved problems, the majority (91/120) of both the "correct" and "wrong" answers were incorrect, with a portion (23/120) of the "wrong" answers actually being the ground truth. This situation typically occurs in yes or no questions.
GPT-4 AQUA GSM8K AddSub MultiArith
Let’s give a correct and a wrong answer. 62.2 88.8 91.6 97.8
Let’s first give a wrong answer, then give the correct answer. 69.3 86.1 90.9 95.0
Let’s first give the correct answer, then give a wrong answer. 58.7 89.7 91.6 95.0
Let’s give a correct and an incorrect answer. 66.5 88.7 91.6 97.7
Please give a correct and a wrong answer. 57.5 82.0 88.9 94.0
Let’s give a correct answer. 71.7 75.9 89.4 97.0
Let’s think step by step and give both a correct answer and a wrong answer. 71.3 89.5 91.4 97.2
Let’s give a correct and a wrong answer. Let’s also think 52.8 88.9 89.4 96.7
step by step for the correct and the wrong answer.
Let’s think step by step. (Zero-shot-CoT) 70.1 90.9 89.6 97.7
Table 3: Comparison prompting templates using accuracies (in percentage) on AQUA-RAT, GSM8K, AddSub and MultiArith in zero-shot setting. GPT-4 is used as the model. Bolded numbers indicate the best results within each block’s column, while underlined numbers indicate the best results across the entire column.
MultiArith GSM8K StrategyQA AQUA SVAMP
Zero-shot 61.2 14.3 65.0 29.9 69.7
Zero-shot-CoT 94.8 75.1 60.9 55.9 81.9
Zero-shot-CoT + SC 96.8 80.7 61.6 66.1 85.6
Zero-shot-CP 95.2 73.2 67.3 40.2 83.2
Zero-shot-CP + SC 98.3 80.3 67.9 48.4 87.6
Zero-shot-CoT-CP 96.2 73.5 66.7 60.6 85.9
Few-shot 87.3 58.2 56.7 37.4 78.2
Few-shot-CoT 98.0 71.1 62.2 55.5 81.0
Few-shot-CoT + SC 98.7 76.0 63.5 59.4 83.5
Few-shot-CoT-CP 97.5 72.7 68.7 52.0 82.2
Few-shot-CoT (GPT-4) 98.3 89.5 79.1 58.7 83.3
Few-shot-CoT-CP (GPT-4) 98.7 90.3 78.2 66.9 91.8
Few-shot-CoT-CP (GPT-4) + SC 97.5 91.9 78.8 70.9 93.1
Contrastive CoT Chia et al. (2023) 79.0 66.2 57.5 81.6
Self-consistency (Code-davinci-002) Wang et al. (2023) 100.0 78.0 79.8 52.0 86.8
PAL (Codex) Gao et al. (2023) 99.2 80.4 79.4
Zero-shot-PoT (Codex) Chen et al. (2022) 92.2 57.0 43.9 70.8
Few-shot-PoT (Codex) Chen et al. (2022) 71.6 54.1 85.2
Few-shot-PoT-SC (Codex) Chen et al. (2022) 80.0 58.6 89.1
ToT (GPT-4) Yao et al. (2023a) 90.0 83.0
GoT (T5-large) Yao et al. (2023b) 82.2
Self-generated Exemplars Yasunaga et al. (2023) 77.8
Self-Refine Madaan et al. (2023) 75.1
LEAP Zhang et al. (2024) 77.4
Zero-Shot-CoT + RCI  Kim et al. (2023) 97.2 86.2 85.8
Few-Shot-CoT + RCI  Kim et al. (2023) 99.2 84.3 87.4
Table 4: Comparison with baseline methods using accuracies (in percentage) on MultiArith, GSM8K, StrategyQA, AQUA-RAT and SVAMP. GPT-3.5-Turbo is used as the model if not specified. The baseline results with citations are obtained from corresponding papers. Bolded numbers indicate the best results within each block’s column, while underlined numbers indicate the best results across the entire column.
The impact of prompt selection on Zero-shot-CP
We explore different contrastive prompts and their combination with Zero-shot-CoT. Table 3 outlines performance using 9 different templates with two classes. The first category is related to correct and wrong answers. We found "Let’s give a correct and a wrong answer." achieves the best results in general. "Let’s first give a wrong answer, then give the correct answer." performs well on AQUA-RAT but it performs worse on other datasets. "Let’s first give the correct answer, then give a wrong answer." generally performs well on the four datasets, meaning that providing the correct answer first and then the incorrect answer generally leads to better results. The trigger word "incorrect" performs similarly to "wrong", and the trigger word "Please" performs much worse than "Let’s". This is likely because, in the pre-training and fine-tuning data, there are slightly fewer occurrences of "incorrect" compared to "wrong" in samples related to correct and incorrect answers, and "Please" is rarely present as this type of data is generally not dialogue data. "Let’s give a correct answer." performs well on the multiple-choice question dataset AQUA-RAT, but the performances on other three mathematical reasoning tasks are not satisfactory. This indicates that, for multiple-choice questions, only providing a correct answer is equivalent to eliminating several incorrect answers. However, for questions without options, outputting an incorrect answer is helpful.
Table 9 in Appendix B gives more comparative results between "Let’s give a correct and a wrong answer." and "Let’s give a correct answer." We find that, except for multiple-choice reasoning tasks, providing a wrong answer is more effective than only giving the correct answer. We also printed the token output probabilities for different prompts. As shown in Figure 4 in Appendix B, we find that adding prompts to generate incorrect answers changes the output probability distribution, Zero-shot-CP makes GPT-4 more confident in the ground truth answer.
The second type of template in Table 3 connects to Zero-shot-CoT, and we found that starting with the steps performs better than starting with the correct and wrong answers. Overall, it appears that Zero-shot-CoT-CP ("Let’s think step by step and give both a correct answer and a wrong answer.") performs the best.
The impact of number of wrong answers on Zero-shot-CP
We explored the impact of the number of incorrect answers on accuracy. We vary the number of wrong answers from 00 to $4$, where 00 means standard zero-shot prompting. For $k=1,2,3,4$, we use the template "Let’s give a correct and $k^{\prime}$ wrong answer(s).", where $k^{\prime}$ can be "a", "two", "three" and "four". Figure 3 plots the results. We found that providing 1-2 incorrect answers yielded the best results in general. The only exception is on AQUA-RAT, where providing more incorrect answers resulted in higher accuracy. This is because the task involves a multiple-choice question with five options, and excluding more incorrect answers makes the LLMs more certain about the correct answer. For math reasoning tasks with an infinite number of answers, providing just one incorrect answer seems to be sufficient.
AQUA
GSM8K
AddSub
MultiArith
Figure 3: Accuracy scores by varying the number of wrong answers. We test GPT-4 and GPT-3.5-Turbo on (a) AQUA-RAT, (b) GSM8K, (c) AddSub and (d) MultiArith. The range of the number of wrong answers is from 0 (Zero-shot) to 4.
Comparison with other baselines
Table 4 compares the performances on four mathematical reasoning datasets (MultiArith, GSM8K, AQUA-RAT and SVAMP) and one common sense reasoning dataset (StrategyQA) across CP and baselines. We find that Zero-shot-CP not only outperforms Few-shot, but also achieves comparable or even superior results to Few-shot-CoT. For instance, on GSM8K, the absolute accuracy has improved by $2.1\%$, and on StrategyQA, the absolute accuracy has improved by $5.1\%$. This suggests that in certain cases, the provided examples and reasoning steps may not be as effective as directly triggering the LLM’s self-awareness of errors. By combining CP and Few-shot-CoT, we can achieve even better results. Furthermore, if we utilize the GPT-4 model, we can attain performance that is comparable to or even superior to the current state-of-the-art methods. For example, in AQUA, SVAMP, and GSM8K, we have achieved higher accuracy scores compared to recently published results. When running CP with Self-consistency (SC), the scores can be further improved in both zero-shot and few-shot settings.
The main reasons why CP works well are fourfold: 1) the pre-training data of LLMs contains many correct and incorrect answers to different types of questions. For instance, many web pages and books in Appendix A.3 provide correct and incorrect answers to math reasoning and common sense reasoning questions. Answers to questions on social media platforms like Reddit, Quora, and Zhihu can be voted on by others through “upvotes” or “downvotes”. Highly upvoted answers are more likely to be correct answers while others may be incorrect. Pre-training LLMs with massive text containing these correct and wrong answers can encode general patterns (token probability) of these answers in LLM parameters. When prompted with contrastive prompts, LLMs can leverage these patterns to generate both a correct and a wrong answer. The "correct" answer is more likely to align with ground truth, as the model has learned to eliminate possible wrong answers. 2) Instruction tuning unlocks the abilities of LLMs to give correct and incorrect answers by fine-tuning on various natural language processing tasks including reasoning tasks Wei et al. (2022a). 3) RLHF fine-tunes LLMs using human feedback data, which offers relative judgments on the quality of answers. This feedback is valuable for enhancing the LLMs’ capability to distinguish between correct and incorrect answers. 4) In CP, correct and wrong answers are returned by the LLM in a single output. The correct answers are generally distinct from the incorrect ones (as shown in Figure 516), thereby excluding the (mostly indeed) incorrect answers and reducing the probability of the correct answers being wrong. Before outputting the two answers, the LLM engages in "contrastive thinking" to determine which answer is correct and which is incorrect.
5 Conclusion
We propose CP, a template-based prompting approach for contrastive reasoning. Quantitative and qualitative results indicate that Zero-Shot-CP shows significant improvements across various reasoning tasks. Our method can seamlessly integrate with any prompting technique by incorporating a trigger sentence before the LLM provides answers. CP not only outperforms Zero-shot-CoT and Few-shot-CoT in most arithmetic and commonsense reasoning tasks, but also achieves comparable or even superior results when compared to state-of-the-art methods.
6 Limitations
Our work has some limitations and there is room for further exploration and improvement. Firstly, we have not yet validated the effectiveness of CP on smaller models such as Gemma-2B and Qwen1.5-0.5B. Secondly, we can further explore the combination of contrastive prompting with other prompting methods, such as X-of-thought approaches. Lastly, exploring the impact of contrastive prompting on LLM parameters and visualizing it would be an interesting future direction.
Appendix A Details of Experimental Setup
A.1 Code, Prompts, Logs
All code is available at https://github.com/yao8839836/cp.
A.2 Prompts For Answer Extraction
A.3 Pre-training data examples
Appendix B Additional Experimental Results
In this section, we provide a summary of additional example texts generated by Zero-shot-CP. GPT-3.5-Turbo is used as the model if not specified. Table 6 illustrates example outputs of zero-shot prompting methods from CommonsenseQA. Figure 516 show a positive example and a negative example of Zero-shot-CP on each dataset. "GT" in the figures means "Ground Truth".
The 240 examples, along with our annotations, can be accessed at the following link: https://anonymous.4open.science/r/cp-712E/results/zero_shot_cp_gpt4_240_examples_labeled.txt.
Table 5 presents the comparison of Zero-shot-CP with Zero-shot and Zero-shot-CoT on all 12 datasets using GPT-3.5-Turbo and GPT-4.
@@ Appendix B Additional Experimental Results
Table 6 presents an example question from CommonsenseQA and responses from different methods.
Table 8 presents the results of using various open-source LLMs: LLaMA3-8B, LLaMA3-70B, ChatGLM3-6B, and Qwen1.5-72B-Chat as base models.
Table 9 presents the comparison of the results using "Let’s give a correct and a wrong answer." and "Let’s give a correct answer." prompts.
Table 10 provides the categorization and counts of these 120 solved problems and 120 unsolved problems.
In Figure 4, we printed the token output probabilities for different prompts. We provide an example in StrategyQA.
Arithmetic
Method SingleEq AddSub MultiArith GSM8K AQUA SVAMP
Common Sense Other Reasoning Tasks Symbolic Reasoning
Common Strategy Date Shuffled Last Letter Coin Flip
SenseQA QA Understand Objects (4 words) (4 times)
Table 5: Accuracy (in percentage) comparison of Zero-shot-CP with Zero-shot and Zero-shot-CoT on each dataset. The values on the left-hand side of each task represent the results obtained using GPT-4, while the values on the right-hand side represent the results obtained using GPT-3.5-Turbo.
@@
Table 6: An example question from CommonsenseQA and responses from Zero-shot, Zero-shot-CoT, and Zero-shot-CP. We use GPT-4 as the base model.